OpenStudy (anonymous):

(Optimization problem, please work through it with me!) Suppose an airline policy states that all baggage must be box-shaped with a sum of length, width, and height not exceeding 108 in. What are the dimensions and volume of a square-based box with the greatest volume under these conditions?

OpenStudy (anonymous):

The constraint is L + W + H = 108 right? And the objective function is V = LWH? I don't know where to go from there...

OpenStudy (anonymous):

Ok so since it is square based that means that I should reformat the constraint to: 2L + H = 108 and the objective function to V = L^2H

OpenStudy (anonymous):

yep, exactly.

OpenStudy (anonymous):

But to be honest, I don't read from the text what they want maximized, minimized, the volume?

OpenStudy (anonymous):

ahh greatest volume!, sorry, it's late here.

OpenStudy (anonymous):

yeah no problem! So next do I solve for L or H? and plug it into the Volume function?

OpenStudy (anonymous):

yes, always make sure that your substitution doesn't make the function complicated, for single variable calculus you want the function to be only dependent on one variable. With simple I mean, always match it to the level of differentiation you are at.

OpenStudy (anonymous):

in this case it doesn't really matter, I am just saying it.

OpenStudy (anonymous):

Ok so in the future I will be able to solve it with multiple variables? But for now the optimization problems will need to be reduced to a single variable?

OpenStudy (anonymous):

yes, they use La Grange multipliers there, don't worry about that just yet. But keep in mind that you always try to express your function in terms of only one variable. For example, this would be such a function: \[ V= f(L)=L^2(108-2L)=108L^2-2L^3 \]

OpenStudy (anonymous):

you see, this function represents the volume, but it only depends on the Length, on nothing else.

OpenStudy (anonymous):

so now I solve for critical points? and how do I know the limit? is it (0, 54) ?

OpenStudy (anonymous):

not limit i mean domain

OpenStudy (anonymous):

\[ f'(x)=0 \] is one condition for it to be an extrema (Maximum Minimum) \[f''(x)<0 \] means that it is a maximum.

OpenStudy (anonymous):

well the domain of the function we have written is \(\mathbb{R}\). But for our solutions, only positive numbers make sense, and zero shouldn't be an answer too.

OpenStudy (anonymous):

\[f'(x)=216L-6L^2=0\] \[ f'(x) = 6L(36-L)=0 \]

OpenStudy (anonymous):

wait so we first find roots and then we find critical points? or do roots not matter?

OpenStudy (anonymous):

excuse me, I am from Switzerland and a lot of the common terms you have don't ring a bell for me. I don't know what you mean with critical points, we actually don't find roots of the function, we find the roots of the derivative which show you where the function takes the highest or lowest point. Here is the graph of the function: http://www.wolframalpha.com/input/?i=108x%5E2-2x%5E3 We are interested in the points where this function has it's highest value (Maximum)

OpenStudy (anonymous):

ohhh ok. Critical point means when slope is zero, what you said

OpenStudy (anonymous):

ahhh, okay, we call that extrema here, or just extreme. But I am glad to know that, thank you (-:

OpenStudy (anonymous):

So i got V'(L) = 216L - 6L^2 when I solved for the derivative

OpenStudy (anonymous):

yup and you want this slope to be equal to zero.

OpenStudy (anonymous):

so when L = 36, the slope is zero

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Now you have basically all the results you need, but make sure, depending on how rigorous your Professor/Teacher is, always make sure that you check with the second derivative what kind of critical point that is.

OpenStudy (anonymous):

In my class they taught us to now plug in this result (L = 36 in our case) to the original function along with the end points of the function (the domain points). And in our case we find the highest value

OpenStudy (anonymous):

but i am not sure what the domain is for this function

OpenStudy (anonymous):

Hmm okay, I think I understand, but to be honest and a bit bold - it makes less sense here to me. Because this Function is defined everywhere, for our purpose we just care about the numbers which are larger than 0 and smaller than infinity.

OpenStudy (anonymous):

So you see, you have nothing to compare it with.

OpenStudy (anonymous):

yes but the sum has to be less than or equal to 108 right?

OpenStudy (anonymous):

Haven't you done second derivative yet to proof that a point is a Maximum or a Minimum? You know that there are two kind of critical points, one is the maximum, where the function takes it's highest values and the one where it's the lowest.

OpenStudy (anonymous):

yeah but what do I compare it with? I know I will have to do f''(36) but what else to make sure?

OpenStudy (anonymous):

yes that's true, the sum has to be less or equal to 108, so you can use that as a constraint, but it's not a real constraint by the function.

OpenStudy (anonymous):

f''(36)<0 you must check that, if that's the case, the point you have found is a maximum and you are done in my opinion.

OpenStudy (anonymous):

ok cool. I guess I just want to be more secure of my answer in case there are other maxima, minima

OpenStudy (anonymous):

In this case, you don't have any other options to check, only if for example your result would have been x=72 72+72=144>108, so this would have been not a correct answer.

OpenStudy (anonymous):

\[f''(x)=216-12L \\ f''(36)=-216<0 \longrightarrow \text{Max}\]

OpenStudy (anonymous):

should read as the second derivative, obviously this latex script isn't working too well in here.

OpenStudy (anonymous):

ok so v(36) is the maximum!

OpenStudy (anonymous):

and I can check again to see if it is a maximum by plugging it into the second derivative, which gives me a negative, which means concave down

OpenStudy (anonymous):

yes it means that it's a maximum, concave down.

OpenStudy (anonymous):

Thank you have a good day! I might have more questions later for other problems.